# Approximating a gene regulatory network from non-sequential data

### v2

@@ Abstract
Given non-sequential snapshots from instances of a dynamical system, we design a compressed sensing based algorithm that reconstructs the dynamical system. We formally prove that successful reconstruction is possible under the assumption that we can construct an approximate clock from a subset of the coordinates of the underlying system.
As an application, we argue that our assumption is likely to be true for RNA-seq datasets, and thus we can recover the underlying nuclear receptor networks and predict pathways, as opposed to genes, that may differentiate between interesting phenotypes in some publicly available datasets.
@@ 1 Introduction
The problem of reconstructing a noisy linear dynamical system from sequential observations to predict its next step is a basic problem in filtering theory. In this paper, we consider the case when the order of observations has been lost, i.e., they are not necessarily sequential. Such a situation arises often in biological samples, where although one can collect information about cells, it is impossible to know perfectly where the cells are in their cell cycle relative to each other. Another area where this problem arises is in reconstructing a linear system learnt over noisy networks, for example in a noisy financial feed where the packets may be disordered or dropped.
Our theoretical contributions: We formulate a concrete version of the above reconstruction problem and solve it under the assumptions discussed in Subsection 1.3. Algorithm 1 in Section 3 is our main algorithmic contribution and Theorem 4.1 in Section 4 upper-bounds the error of the recovered dynamical system using Algorithm 1. The compressed sensing methods used in our paper ensure that reconstruction can be successful, even with small sample sizes, under sparsity assumptions.
Application to genomics: We also show an application of our algorithm to genomic datasets (see Subsection 1.1 for details). We use the reconstructed linear dynamical systems to compute genomic pathways that are prominently different between pathological and reference samples. Thus, unlike standard tools, such as Love et al. (2014), that predict genes whose expressions may differentiate between two phenotypes, we can predict pathways that differentiate between two phenotypes. Figure 1, based on the breast cancer RNA-seq dataset (Cancer Genome Atlas Network: Daniel C Koboldt et al, 2012), represents our main contribution in this context.
(a) Potential pathways for metastasis. The pathological ”phenotype” is presence of metastasis while the reference is absence of metastasis.
(b) Potential pathways which may be adverse to longer survival. The pathological ”phenotype” is survival below 12months while reference is survival beyond 60months.
Figure 1: Predicted pathways where gene A $\to$ gene B means that A (indirectly) up-regulates B, while the flat arrowhead indicates (indirect) down-regulation. A green node indicates that the corresponding gene expression positively correlates with the pathological phenotype (eg. metastasis), while a red node indicates positive correlation with the reference phenotype. Based on the figure, one prominent pathway could be: NR1I2 up-regulates RORB which up-regulates RORC, whose increased expression is positively correlated with metastasis; based on data in Cancer Genome Atlas Network: Daniel C Koboldt et al (2012). See Section 6 for details on how Algorithm 3 was used to recover the pathways from the breast cancer dataset Cancer Genome Atlas Network: Daniel C Koboldt et al (2012).
1.1 Application to genomics
@@
We apply Algorithm 1 to the transcriptional regulation problem in genomics. An important problem in genomics is to understand transcriptional regulation – which genes directly regulate a given gene. For example, one version of the problem is to reconstruct Gene Regulatory Networks (GRNs) from gene expression snapshots measured using RNA-seq data. The surveys Badia-i Mompel et al. (2023) and Barbuti et al. (2020) cover the tools and techniques employed in the study of GRNs. In this paper, we model the dynamical system underlying gene expression as a Markov process. Such models have been used to study gene regulation before (see for example Kepler and Elston (2001) and Chu et al. (2017)). The paper Kepler and Elston (2001) gives a detailed overview of modeling transcription regulation by stochastic differential equations (SDEs); and the Ornstein-Uhlenbeck process that we employ in this paper is just a linear approximation of such SDE models that is analytically tractable and yet expressive enough to capture first order behavior. To the best of our knowledge, we are not aware of prior work that uses $\ell_{1}$ optimization to reconstruct GRNs or the underlying Markov processes.
1.2 Our contribution
Using Algorithms 12 and 3, and the publicly available RNA-seq dataset (Cancer Genome Atlas Network: Daniel C Koboldt et al, 2012), we have predicted pathways, restricted to nuclear receptors (for computational tractability reasons). We show these results in Figure 1 and observe that they may distinguish between metatstatic and non-metastatic breast cancers. See Sections 3.1 and 6 for details.
Note that we do not have the resources to verify the results experimentally, but our predictions may help guide other researchers by narrowing down the genes and pathways for experiments. One way to verify the predictions would be to run our algorithms on a variety of malignant vs control (RNA-seq) datasets and find pathways that are common among the malignant samples. Among the common pathways, one can check for genes that are targets of known drugs. If such drugs exist and are known to be effective, then they help verify the predictions of our algorithms. Conversely, any genes shared amongst common predicted pathways, that are not yet known, can also be used as targets for newer drugs. In other words, our ideas should be applicable to the design of basket trials, where physicians aim to identify and target the same gene or mutation even though they may be treating different cancers. Having a prediction of the underlying pathways may help narrow the number of interesting gene targets that need be explored for therapies. Such trials have been critical in discovering novel therapies like anti-PD1 therapies (for e.g. Le et al. (2017)).
1.3 Underlying assumptions and verification
Our algorithms, specifically Algorithm 1, rely on three assumptions: (1) sparsity: the rows of the matrix, [^1] This is the matrix that constitutes of the coefficients of the first order (degree $1$) terms in the linearization of the underlying dynamical system. are zero (or near zero) in most entries, (2) a mean field model assumption: the expressions in the RNA-seq data-set represent an “average” cell in the tissue concerned, and (3) existence of an approximate clock: we should be able to (roughly) order our cell samples according to cell age (measured as time since last cell division). We now justify these assumptions.
First, the sparsity assumption is expected to hold for GRNs because the transcriptional regulation of a given gene directly relies on only a handful of polymerases, promoters and previous gene products (see, for example Kepler and Elston (2001)). Even if we replace direct regulation by downstream regulation within a few steps, for example when studying GRNs restricted to nuclear receptor genes only, we still expect the resulting matrix to be sparse.
Second, we assume that the population weighted mean expression of any given gene in the tissue sample, that is reflected in the RNA-seq expression value for that gene, is a good approximation of the gene expression in the tissue. Such an assumption is made in mean field theories. The mean field model is usually a valid approximation of a dynamical system when: (a) the sample size is large and (b) fluctuations in value are small (c.f. the Ginzburg criterion). That fluctuations should be small, i.e., the variance for a given gene expression should be small, is intuitive, because biological systems are usually in equilibrium (homeostasis). For an organism to be able to maintain equilibrium, large fluctuations in gene expression should be a rare occurrence, as successfully reverting the system back near its mean is harder to do from far flung outlier values. So if the sample size is large enough, then the mean field assumption should be valid.
Finally, before justifying the approximate clock assumption, we first discuss the reason we assume the existence of an approximate clock. We work with RNA-seq data, which is more readily available, unlike single-cell RNA-seq data. In single-cell RNA-seq data, it is possible to order cells by their latent time (see for example HKU single cell workshop (2021)). However, for RNA-seq data, latent time is not available, as latent time is a model-based construct. One saving grace is that our algorithms do not use the cell sample ordering except to reconstruct the covariance matrix of increments, which is a pretty robust to noise, in that it is only significantly effected if the approximate ordering used has too many errors. Therefore, using appropriate cyclin expression levels to order cell samples in readily available RNA-seq data, although noisy, is sufficiently accurate for our purposes; as evidenced by our verification on a couple of single cell RNA-seq datasets (see Figures 3 and 5).
To elaborate, we verify that the reconstructed covariance matrix of increments is indeed robust to a small number of errors in ordering that may arise from using an approximate clock in three different ways: (1) theoretical justification: we prove that the reconstructed covariance matrix is correct as long as the approximate clock meets certain formal assumptions (Theorem 4.1), (2) by simulation in Section 5: we use synthetic data to study the error in the reconstructed covariance matrix as function of the number of errors in the approximate clock (Figure 2), and (3) by comparing the approximate ordering based on cyclin expressions with the ordering based on cell latent time in murine pancreatic single cell RNA-seq dataset from Bastidas-Ponce et al. (2019) (Figure 3) and murine dentate gyrus dataset of Hochgerner et al. (2018) (Figure 5). In each case, we find that the covariance matrix computed using the approximate clock will be close to the actual covariance matrix.
2 Related problems
@@
Related problems in genomics Our approximate cyclin based ordering problem is related to the problem of trajectory inference, which has been well explored in single cell RNA-seq literature (Saelens, W. and Cannoodt, R. and Todorov, H. et al., 2019). Trajectory inference is likely a harder problem to solve, because we only need to order cell samples well enough to approximately recover a specific covariance matrix in our algorithms, but in the trajectory inference problem one needs to order the cell samples in the correct sequence to predict cellular development.
Finally, it is possible that cyclin expressions themselves may be effected in many situations by the pathology under consideration (see for e.g. Ghafouri-Fard et al. (2022)). In such cases, our algorithms can not be applied directly with cyclins to construct the approximate clock, but it may be possible to use a different set of gene expressions related to cell division, upstream from cyclins, that are verified to be not effected by the pathology, to construct the approximate clock, on a case by case basis.
@@
Related problems in filtering From an abstract algorithmic perspective, the line of work closest to our problem is the Kalman filtering problem in the presence of network loss and delays (see Li et al. (2021)Liu et al. (2020) and Nikfetrat and Esfanjani (2019), and the references therein). However, such results typically assume a independent identical probability of packet drop or delay, sometimes modelled by a Markov process of interest. Also, although the idea of computing conditioned Gaussian distributions (for the error upper-bound in Theorem 4.1) is similar to that used for the analysis of the Kalman filter, the problem studied in our paper differs from the standard filtering problem: (1) we are interested in recovering the underlying linear system and not in predicting the next step of the system, and (2) we use sparsity critically in our algorithms.
2.1 Organization
In Section 3, we formulate the problem and describe our algorithms based on the approximate clock. Next, we verify the existence of an approximate clock is enough to reconstruct the underlying GRN; first theoretically in Section 4 and then empirically in Subsections 5.1 and 5.2. Finally, we give more details on how we obtain the breast cancer pathways in Figure 1 in Section 6.
3 Problem statement and overview
Definition 3.1.
(see for example Oksendal (2013)) Suppose we have a $n$-dimensional Ornstein-Uhlenbeck (OU) process

### v3

@@ Abstract
Given non-sequential snapshots from instances of a dynamical system, we design a compressed sensing based algorithm that reconstructs the dynamical system. On the theoretical side, we show that: (1) successful reconstruction is possible under the assumption that we can construct an approximate clock from a subset of the coordinates of the underlying system, and (2) computing the minimal Lyapunov exponent of the dynamical system, where the minimum is taken over all subsets of coordinates of the dynamical system, equates to computing a min-max equilibrium. We design an efficient randomized algorithm for computing the above equilibrium.
As an application of our theoretical results, we reconstruct the underlying dynamical system from publicly available RNA-seq data to: (1) predict the underlying gene regulatory networks (as opposed to individual genes) that may help differentiate between metastatic vs non-metastatic breast cancer (and also colorectal cancer), and (2) identify candidate genes that could be used as target biomarkers for basket trials. In particular, our in silico analysis suggests that RORC agonists, which are already used in colorectal cancer therapies, may be worth investigating for breast cancers.
@@ 1 Introduction
Reconstructing a noisy linear dynamical system from sequential observations to predict its next step is a basic problem in filtering theory. In this paper, we consider the case when the order of observations has been partially lost. Such a situation arises often in biological samples, where although one can collect (aggregate) information about cells, it is impossible to know perfectly where the cells are in their cell cycle relative to each other. In particular, the relative abundance of bulk RNA-seq data over single cell RNA-seq data in clinical settings was a motivation for formulating the problem. Note that the latter can be temporally ordered in a reasonably error free manner using latent time, but it is not obvious whether the samples in the former can be ordered as well. Our contributions may be viewed from a theoretical lens – abstract algorithms with applications to computational biology, or from a genomic lens – isolating underlying pathological pathways and common biomarkers from various cancers for designing basket trials.
1.1 Theoretical contribution
Consider a linear approximation of a dynamical system: $\dot{x}=Bx+\dot{\eta}$, where $\dot{x}$ denotes the time derivative of $x\in\mathbb{R}^{n}$, $B\in\mathbb{R}^{n\times n}$ is the matrix representing the interactions between the variables, and $\dot{\eta}$ is a white noise term (assumed not too large). If we had a set of ordered samples $x(t),x(t+h),x(t+2h),$... then we could recover the matrix $B$ from these samples by first computing the discrete time derivative to recover $\dot{x}$, and then minimizing an appropriate loss ($\ell_{1}$ loss, for example). If the samples are not ordered, or if we don’t have access to the timestamps $t,t+h,t+2h,..$, then it is not possible to recover the interaction matrix $B$. However, it is natural to ask, if the sample ordering is known up to some small error, for example a small fraction of the ordered samples have been flipped, then can we recover the dynamical system?
We show that one can still recover $B$ by minimizing the $\ell_{1}$ loss. We denote this access to a small error sample ordering as access to an approximate clock. [^1] The terminology comes from our application to genomics, where certain proteins act as approximate clocks. Under the assumption that we have access to an approximate clock, Algorithm 1 in Section 2 allows us to approximately reconstruct the interaction matrix $B$ from the samples. Moreover, we bound the error in approximation of $B$ from Algorithm 1 using Theorems 2.1 and 0.A.1. The error bound in Theorem 0.A.1 is in terms of the amount of error from the actual sample ordering in the approximate clock. Thus we rely on a combination of techniques from stochastic calculus and compressed sensing to show that the second moment of increments of the underlying dynamical system is robust under errors in ordering (see Theorem 0.A.1) and it suffices to reconstruct the underlying sparse dynamical system (Theorem 2.1). This combination of techniques should be novel and interesting. [^2] Although there is some novelty in our modification of the basis pursuit algorithm to enforce row sparsity of matrices (see Theorem 2.1), instead of sparsity of vectors, we do not claim significant theoretical novelty over the proof techniques of dirk in terms of necessary and sufficient criteria for reconstruction of sparse signals. The main utility of compressed sensing methods in Algorithm 1 is to ensure that reconstruction can be successful, even with small sample sizes (relative to underlying dimension, i.e., $n$), which is often the case with RNA-seq data – far fewer patients than number of genes sequenced. Our main application involves reconstructing dynamical systems and predicting pathways responsible for metastasis in breast and colorectal cancers (see Figure 1, and Sections 1.2 and 2.2 for details).
Suppose we successfully reconstruct interaction matrices $B$ and $B^{\prime}$ of two different dynamical systems using Algorithm 1, a natural question is: Do the two given linear dynamical systems evolve differently over a long period of time? More concretely, is there a subset of coordinates $S$ ($S\subset[n]$) such that the orbit corresponding to $B^{\prime}$ projected to $S$ converges to some equilibrium point, but the orbit of $B$ does not converge to any point?
One metric to compare similarity of trajectories is the Lyapunov exponent, which measures whether the long term trajectory of a dynamical system diverges from, or converges to, an equilibrium point. Thus if we find coordinates $S\subset[n]$ such that the Lyapunov exponent corresponding to $B$ (after projection to $S$) is positive and large, but the Lyapunov exponent corresponding to $B^{\prime}$ (after projection to $S$) is significantly smaller than that of $B$ on $S$, then we know that the two dynamical systems have very different long term trajectories on the $S$ coordinates. Since the computation of a Lyapunov exponent for any linear system corresponds to computing the maximum eigenvalue, one needs to compute the minimum (over subsets $S\subseteq[n]$) of the maximum eigenvalue of $B^{\prime}+B^{\prime T}$ after projection; such that the maximum eigenvalue of $B+B^{T}$ is not small, after projection to the coordinates $S$. Two observations are in order:
  1. 1.
    Computing the equilibrium of the min-max objective involves computing a minimum over a potentially exponential number of projections corresponding to subsets of $[n]$ – a non-trivial optimization problem
  2. 2.
    As $|S|$ increases, the maximum eigenvalue of $B^{\prime}+B^{\prime T}$ will be non-decreasing; if $|S|$ decreases, then the maximum eigenvalue of $B+B^{T}$ will be non-increasing. Therefore, the optimal value of $|S|$ may be expected to be neither too small nor too large (see Figure 1(c)).
In Section 2.4, we show that a concrete version of the min-max optimization problem is efficiently solvable by using semidefinite programming (SDP) duality to obtain a SDP relaxation of the problem. We then round the vector solution of the SDP relaxation to a 0-11 integer solution which gives the optimal $S$ and the distance between corresponding Lyapunov exponents (see Algorithm 4). In Theorem 0.B.1, we show that the objective value computed for the SDP relaxation is close to the equilibrium value in expectation. We illustrate one application of the above algorithm: predicting biomarkers for the design of a subclass of clinical trials – basket trials. The idea being to identify common biomarkers between different cancers, so that therapies for one can be translated to another (see Section 2.4 for more details). Currently, such biomarkers are discovered from experiments by trial and error.
1.2 Genomics contribution
In this section, we summarize an application of of our theoretical results above to genomic datasets. Sections 2.22.4 and Supplement 0.A elaborate on our description here. We use the RNA sequencing data on breast cancer metabricmetabric2metabric3, and colorectal cancer rc for our experiments. [^3] About the datasets: The datasets are available at the cbio portal (https://www.cbioportal.org/). The breast cancer dataset contains RNA sequencing data (whole genome from tumor biopsy) from about 2,000 cases, such that half have lymph node involvement (metastasis) and half don’t. The colorectal cancer dataset contains RNA sequencing data (whole genome from tumor biopsy) from about 250 colorectal cancer cases, about 60 of which are non-metastatic (grade T2 or below), while the rest have metastasis (grade T3 or above). Our main contribution, from the perspective of a genomic lens, is an in silico prediction: RORC agonists, which are known to be effective in some colorectal cancers rorc-rx; rorc-rx2, are likely to be effective in some breast cancer cases as well (see also rorc-bc-good; rorc-rx-wbc-ct). More concretely, the steps in the application of our algorithms to obtain the aforementioned conclusion can be summarized as follows:
  1. 1.
    Using Algorithm 1, which assumes the existence of an approximate clock, we reconstruct the noisy linear dynamical system of the underlying gene regulatory network (GRN) for non-metastatic breast and colorectal cancer gene networks (see for e.g., Figures 5(a) and 6(a)), and the corresponding perturbations that lead to metastasis (see for e.g., Figures 5(b), 6(b)). [^4] We restrict our networks to nuclear receptors for computational limitations, but the analysis can be carried for arbitrary sets of genes with bulk RNA sequencing data. The recovered networks are not easily interpretable due to their size (they belong to a $2500$ dimensional space for the $50$ odd nuclear receptors that we consider).
  2. 2.
    We analyze the perturbations (from step (1)), using Algorithm 4, to recover a subset of genes (denoted $\mathcal{C}$), for which the gene expressions for metastatic breast and colorectal cancers will follow similar trajectories, but the evolution of gene expressions of the same set of genes for non-metastatic cases is substantially different for the two cancers (see the discussion in Section 2.4 for details). For breast and colorectal cancers, some of the genes in $\mathcal{C}$ are: RORC, ESR2, NR2F1 (they correspond to the darker areas in the heatmap in Figure 1(a)).
  3. 3.
    Next, we compute the genomic interactions (pathways), using Algorithm 3, that are predicted to be critical for metastasis for breast and colorectal cancers separately. From the computed pathways, we select the top five predicted pathways that contain one or more genes in $\mathcal{C}$ (step (2)). Since $\mathcal{C}$ is supposed to identify genes that behave similarly in breast and colorectal cancer metastasis, one expects to find the common critical interactions from this exercise. Indeed, all five of the resulting pathways for colorectal cancer metastasis were mediated via RORC inhibition, and four of the five pathways for breast cancer metastasis were mediated by RORC inhibition (see Figure 1(c,d)). Thus, we identified RORC inhibition as a common factor in both cancers, even though its exact role may differ for the two cancers. RORC agonists have been investigated for colorectal and other cancer therapies rorc-rx; rorc-rx2. Our in silico results should motivate further investigation of RORC agonists in breast cancer treatment (see also rorc-bc-good; rorc-rx-wbc-ct).
  4. 4.
    Finally, in Section 0.A, we verify that a weighted average of cyclin expressions [^5] These are a family of cellular proteins that regulate the cell cycle, so their concentrations peak at the precise points of time in the cell cycle cccp. For example, Cyclin A is active in the S phase, while Cyclin D is active in the transition from the G1 to S phase. can be used as a sufficiently accurate approximate clock in Algorithm 1. We use single cell data from Bastdas-Ponce and dg, where the latent time from scVelo bergen acts as our ground truth against which we can compare our approximate clock. We show that the cyclin based approximate clock will recover the same covariance matrix, and hence the linear approximation of the gene network, as the latent time based clock (see also Remark 1 and Theorem 0.A.1).
(a) Darker areas identify pairs of genes predicted to have similar expression trajectories in breast and colorectal cancer metastasis. Genes like RORC and NR2F1 could play critical roles in metastasis in both cancers (though their roles may not be the same).
(b) The x-axis shows the dimension $\in[1,50]$, and the y-axis measures the equilibrium value (see Equation 5) corresponding to similarity in trajectories between metastatic breast and colorectal cancers relative to their non-metastatic counterparts. Smaller (negative) values imply higher similarity. Thus, there exist local similarities in gene expressions in metastasis, even though globally breast and colorectal cancers have very different gene expression trajectories (as expected).
(c) Predicted pathways for metastasis in breast cancer. Note that evidence exists that shows up-regulation of NR4A1 slows disease progression nr4a1.
(d) Predicted pathways for metastasis in colorectal cancer. HNF4G expression is known to lead to poor prognosis hnf4g.
Figure 1: Top: The genes in $\mathcal{C}$ obtained using Algorithm 4. See Section 2.4 for details. Bottom: Predicted pathways where gene A $\to$ gene B means that A (indirectly) up-regulates B, while A B indicates (indirect) down-regulation. Based on the Bottom Right figure, one potential pathway for metastasis in colorectal cancer could be: HNF4G down-regulates RORC which down-regulates NR2F2. Thus our algorithm suggests that HNF4G eventually up-regulates NR2F2 via a pathway involving RORC. It is known that increased expression of NR2F2 is associated with metastasis in colorectal cancers nr2f2. See Section 2.2 for details.
Our computations implicitly "average" over inhomogeneous data, and below we discuss resulting limitations.
First, using an approximate clock allows us to overcome the limitation that RNA-seq data is not ordered. However, bulk RNA-seq of tumor specimens contains contributions from many different cell types (cancer cells, stromal cells, immune cells etc). When we reconstruct the approximate GRN (using Algorithm 1), and use it for predicting critical interactions among genes, we are using the averaged expression data in the different cell types in the tumor, as opposed to any specific cell type. Thus our reconstructed model of the GRN does not capture the GRN within any single cell type, and any conclusions drawn from it are not for any particular cell type. Such "averaged" models are often referred to as mean-field approximations, and we discuss them further in Section 0.A.
Second, our in silico results are meant to illustrate the potential of our algorithms . Often the role played by receptors, like RORC, in cancer progression is complex (see for e.g. ror-svy), and using different subcohorts of patients can lead to very different outcomes, due to the inhomogeneous nature of cancer pathways across patients. Using the entire dataset, as we did to obtain Figure 1, can be viewed as "averaging" over many different inhomogeneous pathways from different patient subcohorts. Thus the results obtained may not be true for any given subset of patients. Nevertheless, the fact that we were able to independently predict RORC inhibition as important for disease progression, illustrates that our algorithms may be effective in identifying critical gene targets and pathways, especially if the data is reasonably homogeneous.
1.3 Related work
@@
Modeling GRNs: Understanding transcriptional regulation – which genes directly regulate a given gene – equates to the reconstruction of GRNs from gene expression snapshots. The latter can be measured using RNA-seq data. There is a large volume of work in this area. The surveys grn-svy2,  grn-svy and grn-svy3 cover the tools and techniques employed in the study of GRNs. As in our paper, Markov processes have been used to model gene regulation before (see for example grn1grn2raj, and elowitz). The paper grn1 gives a detailed overview of modeling transcription regulation by stochastic differential equations (SDEs); and the Ornstein-Uhlenbeck process that we employ in this paper is just a linear approximation of such SDE models that is analytically tractable and yet expressive enough to capture first order behavior.
The algorithms for reconstructing GRNs fall under two broad classes based on their input: (1) more recent algorithms work with single cell data as input (for e.g. scenic), while (2) older work (for e.g. arcanegenie) uses gene expression snapshots as the input. Due to the abundance of unordered RNA-seq data vs single cell data in clinical settings, we have chosen to work with the former. However, the results that use unordered RNA-seq data (like clustergraph_model1arcane and genie) do not formally recover the dynamical system; they recover a graphical approximation of the GRN. The crucial difference between the two is that, while there are various ways to define edge weights in any graphical representation (using mutual information arcane, using feature selection genie etc) and that leads to many different ad-hoc similarity measures between such graphs. On the other hand, there is only one unique linear dynamical system approximation of the underlying Markov process for gene regulation. In this paper, we recover an approximation of that representation under the assumptions in Section 0.A. One advantage of recovering the underlying dynamical system is that one can compare similarity in the evolution of trajectories between two such dynamical systems, which corresponds to trajectory of gene expressions. Contrast that with graphical representations, where the walks over such a graph do not necessarily correspond to evolution of trajectories of gene expressions, i.e., the temporal information is not captured.
@@
Relation to trajectory inference: Our approximate cyclin based ordering problem is related to the problem of trajectory inference, which has been well explored in single cell RNA-seq literature clk-svy. Trajectory inference, as opposed to just computing similarity between trajectories via Lyapunov exponents (as in this paper), is likely a harder problem to solve. This is because, we only need to order cell samples well enough to approximately recover a specific covariance matrix in our algorithms, but in the trajectory inference problem one needs to order the cell samples in the correct sequence to predict cellular development. It is possible that cyclin expressions themselves may be effected in many situations by the pathology under consideration (see for e.g. fard). In such cases, our algorithms can not be applied directly with cyclins to construct the approximate clock, but it may be possible to use a different set of gene expressions related to cell division, upstream from cyclins, that are verified to be not effected by the pathology, to construct the approximate clock, on a case by case basis.
Connections to basket trials: A basket trial is a clinical trial aims to group patients by their shared biomarker, as opposed to a disease, and gives them the same treatment that targets the shared biomarker (for e.g. nci-match). Such trials have been critical in discovering novel therapies like anti-PD1 therapies (for e.g. pd1). Algorithm 4 provides a way to predict biomarkers for basket trials. For example, in Figure 1(c,d), we predict that RORC inhibition is common to breast and colorectal cancer metastasis, and RORC agonists have been investigated for colorectal cancers rorc-rx2. This raises the prospect of investigating RORC agonists for breast cancer as well (see also rorc-bc-good). We believe our approach can be useful for predicting biomarker targets across different cancers.
@@
Relation to filtering: Perhaps the line of work closest to Algorithm 1 is Kalman filtering in the presence of network loss and delays (see yangliu and nikfe, and the references therein). However, such results typically assume a independent identical probability of packet drop or delay, sometimes modeled by a Markov process of interest. Also, although the idea of computing conditioned Gaussian distributions (for the error upper-bound in Theorem 0.A.1) is similar to that used for the analysis of the Kalman filter, the problem studied in our paper differs from the standard filtering problem because we are interested in recovering the underlying linear system and not in predicting the next step of the system.
Relation to min-max principles: Perhaps the problem closest to Algorithm 4 (and Equation 5) are versions of the well-known Courant-Fisher min-max principle, which involves a direct computation of the eigenvalues of symmetric matrices by repeatedly computing orthogonal subspaces. However, in our problem, we can not choose the orthogonal subspaces iteratively – the basis is fixed, and we need to compute a single subspace corresponding to a subset of the $[n]$ variables in Equation 5.
2 Technical contributions
In this section, we provide the abstract algorithms (Subsections 2.1 and 2.3), and elaborate on their genomics applications (summarized in Subsection 1.2): step (1) in Subsection 2.2, while steps (2) and (3) in Subsection 2.4. The supplement contains the proofs (Sections 0.A.1 and 0.B), and empirical verification of assumptions (Sections 0.A.2 and 0.A.3).
Definition 1
Suppose we have a $n$-dimensional Ornstein-Uhlenbeck (OU) process